OpenStudy (anonymous):

A truck moves with the velocity of 2 m/s, East and a car moves with the velocity of 4 m/s, East. What is the velocity of the truck with respect to the car? (In other words, if you were sitting in the car, what would you perceive as the velocity of the truck?) A) + 6m/s E B) -6 m/s E C) +2 m/s E D) -2 m/s E

OpenStudy (anonymous):

can you help @random231?

OpenStudy (anonymous):

@bubbles-are-cool.

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

Sorry i have no idea

OpenStudy (anonymous):

That's okay thanks anyway!

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

The velocity of the truck with respect to the car is the velocity of the truck with respect to the ground minus the velocity of the car with respect to the ground. By "minus" I mean you have to use vector subtraction here. \(\color{green}{v_{t/c} = v_{t/g} - v_{c/g} \\ = 2~\text{m/s [E]} - 4~\text{m/s [E]} \\ = 2~\text{m/s [E]} + 4~\text{m/s [W]} \\ = 2~\text{m/s [W]} } \) Imagine that you are sitting in the car looking at the truck. You are advancing further ahead of the truck by 2 metres every second. It appears to you as if the truck is moving backward compared to your position by 2 metres every second.